OpenStudy (hunter68):

Functions f(x) and g(x) are shown below: f(x) = x2 g(x) = x2 + 8x + 16 In which direction and by how many units should f(x) be shifted to obtain g(x)?

OpenStudy (hunter68):

Left by 4 units Right by 4 units Left by 8 units Right by 8 units

OpenStudy (hunter68):

@Hayhayz

OpenStudy (hunter68):

@CandyCove

OpenStudy (seanstewart16):

i got tis mate

OpenStudy (hunter68):

thanks i took this test and got a 20%

OpenStudy (seanstewart16):

those are squared. correct?

OpenStudy (seanstewart16):

not times 2

OpenStudy (hunter68):

yes...

OpenStudy (hunter68):

they are squared...

OpenStudy (seanstewart16):

left four

OpenStudy (hunter68):

i though it was a but it said i was wrong

OpenStudy (seanstewart16):

than your grading system is wrong, or you wrote the equation wrong. I put those into desmos, its not wrong

OpenStudy (hunter68):

i have another one i need help with... do you think you can help?

OpenStudy (seanstewart16):

ya, np

OpenStudy (hunter68):

Functions 1 and 2 are shown below: Function 1: f(x) = −4x2 + 6x + 3 A graph of a parabola with x intercepts negative 1, 0 and 1, 0 and a vertex of 0, 3 is shown. Which function has a larger maximum? Type your answer as 1 or 2.

OpenStudy (hunter68):

i said 1 but i got it wrong...

OpenStudy (seanstewart16):

ill check it

OpenStudy (seanstewart16):

what is function 2?

OpenStudy (hunter68):

okay thanks

OpenStudy (hunter68):

its a graph...

OpenStudy (seanstewart16):

ok

OpenStudy (hunter68):

this is the question...

OpenStudy (hunter68):

#2 this is the exam i need help with

OpenStudy (hunter68):

Was it number 2?

OpenStudy (hunter68):

second time taking this...

OpenStudy (seanstewart16):

number two is wider, therefore, the maximum is greater, yes, it is 2

OpenStudy (hunter68):

ok thanks

OpenStudy (seanstewart16):

np

OpenStudy (hunter68):

now number 3 i said all real numbers and got it wrong...

OpenStudy (hunter68):

Identify the domain of the graph of y = −x2 − 6x − 13.

OpenStudy (hunter68):

All real numbers x ≤ −4 x ≥ −6 x ≥ −2

OpenStudy (hunter68):

thanks for helping me...

OpenStudy (seanstewart16):

np

OpenStudy (hunter68):

i suck at algebra

OpenStudy (seanstewart16):

the domain is the extent of x, so when you graph it, you need to see where it stops, IF it stops

OpenStudy (hunter68):

ok

OpenStudy (seanstewart16):

thats odd

OpenStudy (hunter68):

what?

OpenStudy (seanstewart16):

its a parabola, the domain is infinite

OpenStudy (hunter68):

okay...and that means what?

OpenStudy (seanstewart16):

all real numbers should work

OpenStudy (hunter68):

okay ill try it again

OpenStudy (hunter68):

A soccer ball is kicked toward the goal. The height of the ball is modeled by the function h(t) = −16t2 + 48t where t equals the time in seconds and h(t) represents the height of the ball at time t seconds. What is the axis of symmetry, and what does it represent? t = 1.5; it takes 1.5 seconds to reach the maximum height and 3 seconds to fall back to the ground t = 1.5; it takes 1.5 seconds to reach the maximum height and 1.5 seconds to fall back to the ground t = 3; it takes 3 seconds to reach the maximum height and 3 seconds to fall back to the ground t = 3; it takes 3 seconds to reach the maximum height and 6 seconds to fall back to the ground

OpenStudy (hunter68):

ok this one i said a but got it wrong...

OpenStudy (seanstewart16):

brewski, thats along question. just a minute

OpenStudy (hunter68):

okay

OpenStudy (seanstewart16):

the second one

OpenStudy (hunter68):

okay thanks for checking my answers

OpenStudy (hunter68):

Determine which of the following statements is true concerning the values described in column #1 and column #2. Column #1 Column #2 The x-coordinate of the vertex of the graph of y = −2x2 − 4x + 12 The x-coordinate of the vertex of the graph of y = x2 − 4x + 3 The value found in column #1 is greater than the value found in column #2. The value found in column #1 is less than the value found in column #2. The value found in column #1 is equivalent to the value found in column #2. The relationship between column #1 and column #2 cannot be determined by the information given.

OpenStudy (hunter68):

i said b and got this one wrong

OpenStudy (seanstewart16):

i think a but im not sure

OpenStudy (hunter68):

okay...

OpenStudy (hunter68):

The distance traveled, in feet, of a ball dropped from a tall building is modeled by the equation d(t) = 16t2 where d equals the distance traveled at time t seconds and t equals the time in seconds. What does the average rate of change of d(t) from t = 2 to t = 5 represent? The ball travels an average distance of 112 feet from 2 seconds to 5 seconds. The ball falls down with an average speed of 48 feet per second from 2 seconds to 5 seconds. The ball falls down with an average speed of 112 feet per second from 2 seconds to 5 seconds. The ball travels an average distance of 48 feet from 2 seconds to 5 seconds.

OpenStudy (hunter68):

next one....one more after this...

OpenStudy (hunter68):

D?

OpenStudy (hunter68):

i think...

OpenStudy (hunter68):

Choose the equation below whose axis of symmetry is x = 0. y = x2 + 2x y = x2 − 16x + 58 y = x2 + 2 y = x2 − 4x + 2 last one and i said this was a and got it wrong....

OpenStudy (seanstewart16):

c, and then im workin the last

OpenStudy (hunter68):

thanks for your help...

OpenStudy (seanstewart16):

np

OpenStudy (hunter68):

they are squared...

OpenStudy (hunter68):

x^2*

OpenStudy (seanstewart16):

x^2+2

OpenStudy (hunter68):

submitting it....

OpenStudy (hunter68):

70%

OpenStudy (seanstewart16):

better than 20

OpenStudy (hunter68):

#2 and 5 were t\only the ones that were wrong...

OpenStudy (seanstewart16):

bummer

OpenStudy (seanstewart16):

i gotta go, mssge, if you need anything